The length of an office is x feet and the width is x+2 feet. Write the polynimial that represents the area. Find the area if x = 12.
Answer by jim_thompson5910(35256)

A = LW

A = x(x+2)

A = x^2+2x

A = (12)^2 + 2(12)

A = 144 + 24

A = 168


So the area is 168 square units.